# Convergence order of the quantization error for self-affine measures on Lalley-Gatzouras carpets

### v1

@@ 3.1. Auxiliary coding space
Let $G,G_{y}$ be endowed with discrete topology and let $G^{\mathbb{N}},G_{y}^{\mathbb{N}}$ be endowed with product topology. Then $G^{\mathbb{N}},G_{y}^{\mathbb{N}}$ are both metrizable. The corresponding product metric is compatible with the product topology on $\Phi_{\infty}$. Thus, $\Phi_{\infty}$ is a compact metric space.
With the next lemma, we show that, if $\sigma,\omega\in\Phi^{*}$ and $\sigma]\cap[\omega]\neq\emptyset$, then we have either $\sigma]\subset[\omega]$, or $\omega]\subset[\sigma]$. The proof of this lemma is different from that for Lemma 2.2, because $\Psi^{*},\Phi^{*}$ are endowed with different partial orders.
@@
(1) Assume that $\sigma_{L}\preceq\omega_{L},\omega_{R}\precneqq\sigma_{R}$. For some $p_{1}\geq 0$ and $p_{2}\geq 1$, we write
@@
$\omega=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),\ldots,(\hat{i}_{l+p_{1}},\hat{j}_{l+p_{1}}))\times(j_{l+1},\ldots,j_{k-p_{2}}).$
Then by the definition of $\Phi_{l}$, we have, $b_{\sigma_{y}^{-}}\geq a_{\sigma_{L}}>b_{\sigma_{y}}$. If $p_{1}=0$, we have
@@
This contradicts the fact that $\omega\in]\phi_{l}$. Next, we assume that $p_{1}\geq 1$. Note that $a_{ij}<b_{j}$ for every $(i,j)\in G$. It follows that
@@
This contradicts the fact that $\omega\in\Phi^{*}$. It follows that $|\omega_{R}|\geq|\sigma_{R}|$ and $\sigma_{R}\preceq\omega_{R}$.
@@
    For every $\tau\in G^{l}$, we define $\Omega(\tau):=\{\omega\in G_{y}^{*}:\tau\times\omega\in\Phi_{l}\}$. Then we have
    $\Omega(\tau):=\{\omega\in G_{y}^{*}:b_{\omega^{-}}\geq\frac{a_{\tau}}{b_{\tau_{y}}}>b_{\omega}\}.$
    Hence, $G_{y}^{\mathbb{N}}$ is the disjoint union of the sets $\omega],\omega\in\Omega(\tau)$. Therefore,
    $\bigcup_{\sigma\in\Phi_{l}}[\sigma]=\bigcup_{\tau\in G^{l}}\bigcup_{\omega\in\Omega(\tau)}[\tau\times\omega]=\Phi_{\infty}.$
@@
The following facts have been noted in [30] and will also be useful for the proof of the main theorem.
@@
    $\omega_{L}=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),(i_{l+1},\hat{j}_{l+1})),\sigma_{R}=(j_{l+1},\ldots,j_{k},\ldots,j_{k+p}).$
@@
    $\omega_{L}=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),(i_{l+1},j_{l+1})),\;\sigma_{R}=(j_{l+2},\ldots,j_{k},\ldots,j_{k+p}).$
@@ 3.2. Auxiliary measures
For $h\geq 1,l\geq 1$ and $\sigma\in\Phi_{l}$, we define
$\mathcal{E}_{r}(\sigma):=\mathcal{E}_{r}(\mathcal{L}^{-1}(\sigma)),\;\Lambda_{h}(\sigma):=\{\rho\in\Phi_{l+h}:\sigma\preceq\rho\};$
$I_{h,r}(t):=\sum_{\omega\in\Phi_{h}}\mathcal{E}_{r}(\omega)^{t},\;t>0.$
@@
We divide the proof of (3.1) into the subsequent three lemmas.
@@
Let $f_{1}(t):=\min\{1,(\sum_{j\in G_{y}}q_{j}^{t})^{-1}\}$. For $\sigma\in\Phi^{*}$ and $h\geq 1$, we have
$\sum_{\rho\in\Lambda_{h}(\sigma)}\mathcal{E}_{r}(\rho)^{t}\geq f_{1}(t)\mathcal{E}_{r}(\sigma)^{t}I_{h,r}(t).$
Proof.
[…]
It follows that
$\prod_{p=1}^{k-1}b_{j_{p}}\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}\geq\prod_{p=1}^{l}a_{i_{p}j_{p}}\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}>\prod_{p=1}^{k}b_{j_{p}}\prod_{p=1}^{\hat{k}}b_{j_{p}}.$
We need to distinguish between the following two cases:
  1. (1)
    $\prod_{p=1}^{k}b_{j_{p}}\prod_{p=1}^{\hat{k}-1}b_{j_{p}}\geq\prod_{p=1}^{l}a_{i_{p}j_{p}}\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}$. In this case, we define
    We have that $\rho(\omega)\in\Lambda_{h}(\sigma)$. We denote by $\Lambda_{h,1}(\sigma)$ the of set such words $\rho(\omega)$ and denote the set of the words $\omega$ by $\Phi_{h,1}$.
  2. (2)
    $\prod_{p=1}^{k}b_{j_{p}}\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}<\prod_{p=1}^{l}a_{i_{p}j_{p}}\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}$. We define
    $\rho(\omega):=(\sigma_{L}\ast\omega_{L})\times(j_{l+1},\ldots,j_{k-1},\hat{j}_{h+1},\ldots,\hat{j}_{\hat{k}-1},j_{k}).$
    Then $\rho(\omega)\in\Lambda_{h}(\sigma)$. We denote the set of such words $\rho(\omega)$ by $\Lambda_{h,2}(\sigma)$ and denote the set of the words $\omega$ by $\Phi_{h,2}$.
Let $\tau\in G_{y}^{*}$. If $|\tau|=1$, we define $\tau^{\sharp}:=\theta$; otherwise, we denote by $\tau^{\sharp}$ the word that is obtained by deleting the first letter of $\tau$. We observe the following facts.
  1. (i)
    We have, $\Phi_{h,1}\cap\Phi_{h,2}=\emptyset,\Lambda_{h,2}(\sigma)\cap\Lambda_{h,2}(\sigma)=\emptyset$. The first equality is clear, so we need only to show the second. In fact, a word $\rho(\omega)\in\Lambda_{h,1}(\sigma)$ can not be obtained by any $\tau\in\Phi_{h,2}$ and vice versa. Otherwise, we would have $\omega_{L}=\tau_{L},\tau_{R}|_{1}=j_{k}$ and $\omega_{R}=(\tau_{R}^{-})^{\sharp}\ast j_{k}$. This leads to
    $b_{\omega_{y}}=b_{(\omega_{L})_{y}}b_{\omega_{R}}=b_{(\tau_{L})_{y}}b_{\omega_{R}}=b_{(\tau_{L})_{y}}b_{(\tau_{R}^{-})^{\sharp}\ast j_{k}}=b_{\tau^{-}}\geq a_{\tau_{L}}=a_{\omega_{L}},$
    contradicting the fact that $\omega\in\Phi_{h}$.
  2. (ii)
    For different words $\omega,\tau\in\Phi_{h,1}$, we have $\rho(\omega)\neq\rho(\tau)$.
  3. (iii)
    There exist at most $m$ words in $\Phi_{h,2}$ that determine the same word in $\Lambda_{h,2}(\sigma)$, because of the absence of $\hat{j}_{\hat{k}}$ in $\rho(\omega)$. Fix an $\omega\in\Phi_{h,2}$. For $j\in G_{y}$, let $\hat{\omega}^{(j)}:=\omega_{L}\times(\omega_{R}^{-}\ast j)$. Whenever $\hat{\omega}^{(j)}\in\Phi_{h}$, we have, $\rho(\hat{\omega}^{(j)})=\rho(\omega)$ and $\hat{\omega}^{(j)}\in\Phi_{h,2}$. Hence, we obtain that (cf. Remark 3.5 below)
    $\rho^{-1}(\rho(\omega))=\{\hat{\omega}^{(j)}:j\in G_{y}\}\cap\Phi_{h}.$
    We write $\langle\omega\rangle:=\{\hat{\omega}:\rho(\hat{\omega})=\rho(\omega)\}$. For every $\omega\in\Phi_{h}(2)$, we take an arbitrary word of $\langle\omega\rangle$ and denote the set of all such words by $\Phi^{\flat}_{h}(2)$.
Using the above facts, we deduce that
$\sum_{\rho\in\Lambda_{h}(\sigma)}\mathcal{E}_{r}(\rho)^{t}$ $\geq$ $\sum_{\rho\in\Lambda_{h,1}(\sigma)}\mathcal{E}_{r}(\rho)^{t}+\sum_{\rho\in\Lambda_{h,2}(\sigma)}\mathcal{E}_{r}(\rho)^{t}$
$=$ $\mathcal{E}_{r}(\sigma)^{t}\sum_{\omega\in\Phi_{h}(1)}\mathcal{E}_{r}(\omega)^{t}+\sum_{\omega\in\Phi^{\flat}_{h}(2)}\mathcal{E}_{r}(\rho(\omega))^{t}$
$=$ $\mathcal{E}_{r}(\sigma)^{t}\sum_{\omega\in\Phi_{h}(1)}\mathcal{E}_{r}(\omega)^{t}+\mathcal{E}_{r}(\sigma)^{t}\sum_{\omega\in\Phi^{\flat}_{h}(2)}(p_{\omega_{L}}q_{\omega_{R}^{-}}a_{\omega_{L}}^{r})^{t}$
$\geq$ $\mathcal{E}_{r}(\sigma)^{t}\sum_{\omega\in\Phi_{h}(1)}\mathcal{E}_{r}(\omega)^{t}+\mathcal{E}_{r}(\sigma)^{t}\sum_{\omega\in\Phi_{h}(2)}\mathcal{E}_{r}(\omega)^{t}\big{(}\sum_{j\in G_{y}}q_{j}^{t}\big{)}^{-1}$
$\geq$ $f_{1}(t)\mathcal{E}_{r}(\sigma)^{t}\sum_{\omega\in\Phi_{h}}\mathcal{E}_{r}(\omega)^{t}.$
This completes the proof of the lemma. ∎
Remark 3.5.
Assume that $\omega\in\Phi_{h,2}$ and for some $j\in G_{y}$, $\hat{\omega}^{(j)}\notin\Phi_{h}$. Then there exists some $p\geq 1$ such that $\hat{\omega}^{(j)+}:=\omega_{L}\times(\omega_{R}^{-}\ast(j,\widetilde{j}_{1},\ldots,\widetilde{j}_{p}))\in\Phi_{h,2}$. To see this, let $\sigma,\omega$ be the same as in Lemma 3.4. Then (3.2) holds. Since $\hat{\omega}^{(j)}\notin\Phi_{h}$, we have, $\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}b_{j}>\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}$. Thus, there exists some integer $p\geq 1$, such that
$\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}\cdot b_{j\ast\widetilde{j}_{1},\ldots,\widetilde{j}_{p-1}}\geq\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}>\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}\cdot b_{j\ast\widetilde{j}_{1},\ldots,\widetilde{j}_{p}}.$
Since $\omega\in\Phi_{h,2}$, we deduce that
$\prod_{p=1}^{k}b_{j_{p}}\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}\cdot b_{j\ast\hat{j}_{1}\cdots\ast\hat{j}_{p-1}}<\prod_{p=1}^{k}b_{j_{p}}\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}<\prod_{p=1}^{l}a_{i_{p}j_{p}}\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}.$
This implies that $\omega_{L}\times(\omega_{R}^{-}\ast(j,\widetilde{j}_{1},\ldots,\widetilde{j}_{p}))\in\Phi_{h,2}$. Note that
$\prod_{p=1}^{k}b_{j_{p}}\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}<\prod_{p=1}^{l}a_{i_{p}j_{p}}\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}\leq\prod_{p=1}^{k-1}b_{j_{p}}\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}\cdot b_{j}.$
We clearly have $b_{j}>b_{\hat{j}_{\hat{k}}}$. We have
$\rho(\hat{\omega}^{(j)+})=(\sigma_{L}\ast\omega_{L})\times(\sigma_{R}\ast\omega_{R}^{-}\ast(j,\widetilde{j}_{1},\ldots,\widetilde{j}_{p-1},j_{k}).$
Analogously to Lemma 2.4, we have
Lemma 3.6.
Let $A_{6}:=[A_{4}^{-1}]$ and $\sigma\in\Phi^{*}$. Then (1) for every $h>A_{6}$ and $\rho\in\Lambda_{h}(\sigma)$, we have $|\rho_{R}|\geq|\sigma_{R}|+1$; (2) for every $\rho\in\Lambda_{1}(\sigma)$, we have $|\rho_{R}|-|\sigma_{R}|\leq A_{1}$.
@@
$\prod_{p=1}^{l+h}a_{i_{h}j_{h}}$ $=$ $\prod_{p=1}^{l}a_{i_{h}j_{h}}\prod_{p=1}^{h}a_{\hat{i}_{h}\hat{j}_{h}}\leq\prod_{p=1}^{k-1}b_{j_{h}}\prod_{p=1}^{h}a_{i_{h}j_{h}}$
$\leq$ $\underline{b}^{-1}\prod_{p=1}^{k}b_{j_{h}}\prod_{p=1}^{h}a_{\hat{i}_{h}\hat{j}_{h}}=\underline{b}^{-1}\prod_{p=1}^{k}b_{j_{h}}\prod_{p=1}^{h}b_{\hat{j}_{p}}\prod_{p=1}^{h}\frac{a_{\hat{i}_{h}\hat{j}_{h}}}{b_{\hat{j}_{p}}}$
$\leq$ $\underline{b}^{-1}A_{3}^{h}\prod_{p=1}^{k}b_{j_{h}}\prod_{p=1}^{h}b_{\hat{j}_{p}}$
$<$ $\prod_{p=1}^{k}b_{j_{h}}\prod_{p=1}^{h}b_{\hat{j}_{p}}$
@@
Remark 3.7.
For $\sigma\in\Phi^{*}$ and $h\geq 1$, we have ${\rm card}(\Lambda_{h}(\sigma))\leq N^{h}m^{hA_{1}}$, where $N:={\rm card}(G)$. Therefore,
$\sum_{\rho\in\Lambda_{h}(\sigma)}\mathcal{E}_{r}(\rho)^{t}\leq N^{h}m^{hA_{1}}\overline{\eta}_{r}^{ht}\mathcal{E}_{r}(\sigma)^{t}.$
Let $\xi(t):=\max\limits_{1\leq h\leq A_{6}}(I_{h,r}(t)^{-1}N^{h}m^{hA_{1}}\overline{\eta}_{r}^{ht})$. Then for every $1\leq h\leq A_{6}$, we have
$\sum_{\rho\in\Lambda_{h}(\sigma)}\mathcal{E}_{r}(\rho)^{t}\leq\xi(t)\mathcal{E}_{r}(\sigma)^{t}I_{h,r}(t).$
@@
Let $f_{2}(t):=\max\{\xi(t),\underline{q}^{-t}\}$. For every $\sigma\in\Phi_{l}$ and $h\geq 1$, we have
(3.3) $\sum_{\rho\in\Lambda_{h}(\sigma)}\mathcal{E}_{r}(\rho)^{t}\leq f_{2}(t)\mathcal{E}_{r}(\sigma)^{t}I_{h,r}(t).$
@@
Let $\sigma=((i_{1},j_{1}),\ldots,(i_{l},j_{l}))\times(j_{l+1},\ldots,j_{k})\in\Phi_{l}$. By Remark 3.7, it suffices to prove (3.3) for $h\geq A_{6}$. Let $\rho$ be an arbitrary word in $\Lambda_{h}(\sigma)$. We write
$\rho_{L}=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),(\hat{i}_{1},\hat{j}_{1}),\ldots,(\hat{i}_{h},\hat{j}_{h})),$
@@
$\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}>\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}>b_{j_{k}}\prod_{p=1}^{\hat{k}}b_{j_{p}}.$
@@
    $\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}>\prod_{p=1}^{\hat{k}}b_{j_{p}}$. In this case, we define
@@
    Then $\omega(\rho)\in\Phi_{h}$. Let $\Phi_{h,1}$ denote the set of all such words $\omega(\rho)$ and let $\Lambda_{h,1}$ denote the set of the words $\rho$ in this case.
@@
    $\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}\leq\prod_{p=1}^{\hat{k}}b_{j_{p}}$. In this case, we define
@@
    One can see that $\omega(\rho)\in\Phi_{h}$. Let $\Phi_{h,2}$ denote the set of all such words $\omega$ and let $\Lambda_{h,2}$ the set of the words $\rho$ in this case.
We clearly have that $\Lambda_{h,1}\cap\Lambda_{h,2}=\emptyset$. Also, we have $\Phi_{h,1}\cap\Phi_{h,2}=\emptyset$. Otherwise, there would exist some $\rho^{(1)}\in\Lambda_{h,1}$ and $\rho^{(2)}\in\Lambda_{h,2}$ such that $\omega:=\omega(\rho^{(1)})=\omega(\rho^{(2)})=:\tau$. Then we have, $\omega_{L}=\tau_{L}$ and $\tau_{R}=\omega_{R}^{-}$ contradicting (3.4). It follows that
$I_{h,r}(t)$ $\geq$ $\sum_{\omega\in\Phi_{h,1}}\mathcal{E}_{r}(\omega)^{t}+\sum_{\omega\in\Phi_{h,2}}\mathcal{E}_{r}(\omega)^{t}$
$\geq$ $\sum_{\rho\in\Lambda_{h,1}}\frac{\mathcal{E}_{r}(\rho)^{t}}{\mathcal{E}_{r}(\sigma)^{t}}+\underline{q}^{t}\sum_{\rho\in\Lambda_{h,2}}\frac{\mathcal{E}_{r}(\rho)^{t}}{\mathcal{E}_{r}(\sigma)^{t}}$
@@
This completes the proof of the lemma. ∎
Combining Lemmas 3.4, 3.8, we obtain the next lemma, which will be used in the construction of auxiliary measures.
@@
For $k,p\geq 1$, we have
$f_{1}(t)I_{k,r}(t)I_{p,r}(t)\leq I_{k+p,r}(t)\leq f_{2}(t)I_{k,r}(t)I_{p,r}(t).$

### v10

@@ 3.1. Auxiliary coding space
Let $G,G_{y}$ be endowed with discrete topology and let $G^{\mathbb{N}},G_{y}^{\mathbb{N}}$ be endowed with product topology. Then $G^{\mathbb{N}},G_{y}^{\mathbb{N}}$ are both metrizable. The corresponding product metric is compatible with the product topology on $\Phi_{\infty}=G^{\mathbb{N}}\times G_{y}^{\mathbb{N}}$. Thus, $\Phi_{\infty}$ is a compact metric space.
With the next lemma, we show that, if $\sigma,\omega\in\Phi^{*}$ and $\sigma]\cap[\omega]\neq\emptyset$, we have either $\sigma]\subset[\omega]$, or $\omega]\subset[\sigma]$. The proof of this lemma is different from that for Lemma 2.1, because $\Psi^{*},\Phi^{*}$ are endowed with different partial orders.
@@
(1) Assume that $\sigma_{L}\preceq\omega_{L},\omega_{R}\precneqq\sigma_{R}$. We write
@@
$\omega=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),\ldots,(\hat{i}_{l+p_{1}},\hat{j}_{l+p_{1}}))\times(j_{l+1},\ldots,j_{k-p_{2}}),$
where $p_{1}\geq 0$ and $p_{2}\geq 1$. We have, $b_{\sigma_{y}^{-}}\geq a_{\sigma_{L}}>b_{\sigma_{y}}$. If $p_{1}=0$, we have
@@
This contradicts the fact that $\omega\in\Phi_{l}$. Next, we assume that $p_{1}\geq 1$. Note that $a_{ij}<b_{j}$ for every $(i,j)\in G$. It follows that
@@
This contradicts the fact that $\omega\in\Phi_{l+p_{1}}$. Hence, $|\omega_{R}|\geq|\sigma_{R}|$ and $\sigma_{R}\preceq\omega_{R}$.
@@
  1. (r0)
    Let $\sigma=((i_{1},j_{1}),\ldots,(i_{l},j_{l}))\times(j_{l+1},\ldots,j_{k})\in\Phi_{l}$ and $\omega\in\Phi_{l+1}$ with $\sigma\preceq\omega$. Then for some $p\geq 0$ and $(i,j)\in G$, we have
    $\omega=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),(i,j))\times(j_{l+1},\ldots,j_{k},\ldots,j_{k+p}).$
@@
    For every $\omega\in G^{l}$, we define $\Omega(\omega):=\{\tau\in G_{y}^{*}:\omega\times\tau\in\Phi_{l}\}$. We have
    $\Omega(\omega):=\{\tau\in G_{y}^{*}:b_{\tau^{-}}\geq\frac{a_{\omega}}{b_{\omega_{y}}}>b_{\tau}\}.$
    Hence, $G_{y}^{\mathbb{N}}$ is the disjoint union of the sets $\tau],\tau\in\Omega(\omega)$. Therefore,
    $\bigcup_{\sigma\in\Phi_{l}}[\sigma]=\bigcup_{\omega\in G^{l}}\bigcup_{\tau\in\Omega(\omega)}[\omega\times\tau]=\Phi_{\infty}.$
@@
The following facts will also be useful (cf. [26]).
@@
    $\omega_{L}=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),(\hat{i}_{l+1},\hat{j}_{l+1})),\;\omega_{R}=(j_{l+1},\ldots,j_{k},\ldots,j_{k+p}).$
@@
    $\omega_{L}=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),(i_{l+1},j_{l+1})),\;\omega_{R}=(j_{l+2},\ldots,j_{k},\ldots,j_{k+p}).$
@@ 3.2. Auxiliary measures
For $\sigma\in\Psi^{*}$, let $\mathcal{E}_{r}(\sigma)$ be as defined in (2.2). We define
$\mathcal{E}_{r}(\sigma):=\mathcal{E}_{r}(\mathcal{L}^{-1}(\sigma)),\;\sigma\in\Phi^{*};$
$\Lambda_{h}(\sigma):=\{\rho\in\Phi_{l+h}:\sigma\preceq\rho\};\;I_{h,r}(t):=\sum_{\omega\in\Phi_{h}}\mathcal{E}_{r}(\omega)^{t},\;t\geq 0,\;h\geq 1.$
@@
For $t=0$, (3.1) trivially becomes ${\rm card}(\Lambda_{h}(\sigma))\asymp{\rm card}(\Phi_{h})$. In the following, we divide the proof of (3.1) into three lemmas.
@@
Let $A_{6}:=[A_{4}^{-1}]$ and $\sigma\in\Phi^{*}$. (1) for every $h>A_{6}$ and $\rho\in\Lambda_{h}(\sigma)$, we have $|\rho_{R}|\geq|\sigma_{R}|+1$; (2) for every $\rho\in\Lambda_{1}(\sigma)$, we have $|\rho_{R}|-|\sigma_{R}|\leq A_{1}$.
@@
$a_{\rho_{L}}$ $=$ $\prod_{p=1}^{l}a_{i_{p}j_{p}}\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}\leq\prod_{p=1}^{k-1}b_{j_{p}}\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}$
$\leq$ $\underline{b}^{-1}\prod_{p=1}^{k}b_{j_{p}}\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}=\underline{b}^{-1}\prod_{p=1}^{k}b_{j_{p}}\prod_{p=1}^{h}b_{\hat{j}_{p}}\prod_{p=1}^{h}\frac{a_{\hat{i}_{p}\hat{j}_{p}}}{b_{\hat{j}_{p}}}$
[…]
$<$ $\prod_{p=1}^{k}b_{j_{p}}\prod_{p=1}^{h}b_{\hat{j}_{p}}$
@@
For every $t\geq 0$, there exists a number $h_{1,r}(t)>0$ such that for every $\sigma\in\Phi^{*}$ and $h\geq 1$, the following holds:
(3.2) $\sum_{\rho\in\Lambda_{h}(\sigma)}\mathcal{E}_{r}(\rho)^{t}\leq h_{1,r}(t)\mathcal{E}_{r}(\sigma)^{t}I_{h,r}(t).$
@@
Let $N:={\rm card}(G)$. For $\sigma\in\Phi^{*}$ and $h\geq 1$, by Lemma 3.4,, we have ${\rm card}(\Lambda_{h}(\sigma))\leq N^{h}m^{hA_{1}}$. Hence,
$\sum_{\rho\in\Lambda_{h}(\sigma)}\mathcal{E}_{r}(\rho)^{t}\leq N^{h}m^{hA_{1}}(\overline{p}\overline{a}^{r})^{h}\mathcal{E}_{r}(\sigma)^{t}<N^{h}m^{hA_{1}}\overline{\eta}_{r}^{ht}\mathcal{E}_{r}(\sigma)^{t}.$
Let $\xi_{1,r}(t):=\max\limits_{1\leq h\leq A_{6}}(I_{h,r}(t)^{-1}N^{h}m^{hA_{1}}\overline{\eta}_{r}^{ht})$. For every $1\leq h\leq A_{6}$, we have
$\sum_{\rho\in\Lambda_{h}(\sigma)}\mathcal{E}_{r}(\rho)^{t}\leq\xi_{1,r}(t)\mathcal{E}_{r}(\sigma)^{t}I_{h,r}(t).$
Next, we assume that $h>A_{6}$. Let $\sigma=((i_{1},j_{1}),\ldots,(i_{l},j_{l}))\times(j_{l+1},\ldots,j_{k})$ be given. Let $\rho$ be an arbitrary word in $\Lambda_{h}(\sigma)$. By Lemma 3.4, we have $|\rho_{R}|\geq|\sigma_{R}|+1$. Write
@@
$\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}>\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}>b_{j_{k}}\prod_{p=1}^{\hat{k}}b_{\hat{j}_{p}}.$
@@
    $\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}>\prod_{p=1}^{\hat{k}}b_{\hat{j}_{p}}$. In this case, we define
@@
    Then $\omega(\rho)\in\Phi_{h}$. Let $\widetilde{\Phi}_{h,1}$ denote the set of all such words $\omega(\rho)$ and let $\widetilde{\Lambda}_{h,1}(\sigma)$ denote the set of the words $\rho$ in this case.
@@
    $\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}\leq\prod_{p=1}^{\hat{k}}b_{\hat{j}_{p}}$. In this case, we define
@@
    One can see that $\omega(\rho)\in\Phi_{h}$. Let $\widetilde{\Phi}_{h,2}$ denote the set of all such words $\omega$ and let $\widetilde{\Lambda}_{h,2}(\sigma)$ the set of the words $\rho$ in this case.
We clearly have that, $\widetilde{\Lambda}_{h,1}(\sigma)\cap\widetilde{\Lambda}_{h,2}(\sigma)=\emptyset,\Lambda_{h}(\sigma)=\widetilde{\Lambda}_{h,1}(\sigma)\cup\widetilde{\Lambda}_{h,2}(\sigma)$. Further, we have that $\widetilde{\Phi}_{h,1}\cap\widetilde{\Phi}_{h,2}=\emptyset$. Otherwise, there would exist some $\rho^{(1)}\in\widetilde{\Lambda}_{h,1}(\sigma)$ and $\rho^{(2)}\in\widetilde{\Lambda}_{h,2}(\sigma)$ such that $\omega(\rho^{(1)})=\omega(\rho^{(2)})$. Then $(\rho^{(1)})_{L}=(\rho^{(2)})_{L}$ and $(\rho^{(2)})_{R}=((\rho^{(1)})_{R})^{-}$, contradicting (3.3). For $t=0$, we simply have that ${\rm card}(\Lambda_{h}(\sigma))\leq{\rm card}(\Phi_{h})$. For $t>0$, we have
$I_{h,r}(t)$ $\geq$ $\sum_{\omega\in\widetilde{\Phi}_{h,1}}\mathcal{E}_{r}(\omega)^{t}+\sum_{\omega\in\widetilde{\Phi}_{h,2}}\mathcal{E}_{r}(\omega)^{t}$
$\geq$ $\sum_{\rho\in\widetilde{\Lambda}_{h,1}(\sigma)}\frac{\mathcal{E}_{r}(\rho)^{t}}{\mathcal{E}_{r}(\sigma)^{t}}+\underline{q}^{t}\sum_{\rho\in\widetilde{\Lambda}_{h,2}(\sigma)}\frac{\mathcal{E}_{r}(\rho)^{t}}{\mathcal{E}_{r}(\sigma)^{t}}$
@@
It is sufficient to define $h_{1,r}(t):=\max\{\xi_{1,r}(t),\underline{q}^{-t}\}$. ∎
@@
For every $t\geq 0$, there exists a number $h_{2,r}(t)>0$, such that for every $\sigma\in\Phi^{*}$ and $h\geq 1$, the following holds:
(3.4) $\sum_{\rho\in\Lambda_{h}(\sigma)}\mathcal{E}_{r}(\rho)^{t}\geq h_{2,r}(t)\mathcal{E}_{r}(\sigma)^{t}I_{h,r}(t).$